If you purchased a U.S. Treasury at a price of $105:08 with a par amount of $5,000, how much did
you pay? _______
A) $5,000
B) $5,250
C) $4,762
D) $5,262
Answer #1

Fraction of 08 = 08/32 = 25%

Price paid is at 105.25%

Payment for the bond = 5000 * 105.25 % = 5262.5

Ans : 5262 (rounded off)